OpenStudy (anonymous):

Solve 4(c - 3) = 4c + 8. A. c = 11 B. c = 5 C. c = 0 D. No Solution

OpenStudy (anonymous):

4c-12=4c+8 4c-4c=8+12 0=20 ----->No solution (D)

OpenStudy (anonymous):

it was no solution @eyad is right^^